# Black-Hole Cartography

### v1

@@ IV Quadratic QNM results
Second-order PT does not give a clear, unique prediction for the spatial shape of second-order QQNMs. Because the QQNM is sourced by two linear QNMs, it is expected that the shape $F_{\alpha\alpha^{\prime}}$ (see Eq. 5) will be given by some quadratic combination of spheroidal harmonics.
Prediction A: A naive guess for these functions is to take the product of two parent spheroidal harmonics: ${}_{-2}S_{\alpha}(\theta,\phi)$ and ${}_{-2}S_{\alpha^{\prime}}(\theta,\phi)$. Then the spatial shape of the QQNM takes the form
@@
However, while this gives qualitatively the correct dependence on the azimuthal $\phi$ coordinate, it gives the wrong $\theta$-dependence; see, for example, the reconstruction of the dominant QQNM in Fig. 7. In particular, it predicts that the dominant QQNM’s largest observed amplitude would be at the north pole ($\theta=0$). This is in disagreement with the reconstruction, which shows a finite angle $\theta_{\rm max}\approx 35^{\circ}$ for this maximum. This is further quantified in Fig. 8, where $\theta_{\mathrm{max}}$ is plotted for the $\alpha=\alpha^{\prime}=(2,2,0,+)$ QQNM for all simulations.
This prediction cannot be correct because the resulting $F_{\alpha\alpha^{\prime}}$ does not have the required -2 spin weight for a GW field. Instead $F_{\alpha\alpha^{\prime}}$ in Eq. 18 has spin-weight -4 resulting from the fact that the product of a spin-weight $s$ function and a spin-weight $s^{\prime}$ function has spin-weight $s+s^{\prime}$.
Prediction B: This issue can be resolved by using a product of spheroidal harmonics with spin weights which sum to -2. There is not a unique choice for this. The possibilities for the fundamental QQNM with $\alpha=\alpha^{\prime}=(2,2,0,+)$ are
@@
Prediction C: There is another proposal given in [42]. This assumes that the shape of the QQNM is given by a single spheroidal harmonic which, rather than being defined by its indices, has a modified spherodicity defined using the frequency $\omega_{\alpha}+\omega_{\alpha^{\prime}}$ of the QQNM,
$\mathbf{C:}\quad F_{\alpha\alpha^{\prime}}(\theta,\phi)\propto{}_{-2}S_{\lambda\mu}\big{(}\theta,\phi;\chi_{f}\left[\omega_{\alpha}+\omega_{\alpha^{\prime}}\right]\big{)},$ (21)
where $\lambda=\ell+\ell^{\prime}$ and $\mu=m+m^{\prime}$. Here, the notation for the spheroidal harmonics from Ref. [43] is used.
Prediction D: An alternative resolution to the spin-weight issue is to take the product of two parent spheroidal harmonics with spin weights which do not sum to -2, but act on them with the appropriate spin-weight raising or lowering operator to obtain the required spin-weight -2 function. For example, taking the product of two spin-weight -2 spheroidal harmonics as in prediction A but acting on them twice using the spin-weight raising operator gives,
@@
For each of these predictions, a corresponding quadratic mode mixing coefficient $\mu^{\ell m}_{\alpha\alpha^{\prime}}$ can be computed. These three predictions are compared in Fig. 7 against the reconstruction of the fundamental QQNM in simulation 0001. Prediction A is clearly disfavoured and is not discussed further. Visually, there is little difference between the other predictions.
Interestingly, the B, C and D predictions all correctly suggest that the projected amplitude of the fundamental QQNM is maximized at a non-zero viewing angle, away from the axis of the remnant BH’s final spin. This has important consequences for the observability of this QQNM in GW data. This angle is investigated further in Fig. 8 where it is plotted as a function of the spin magnitude of the remnant for all remaining predictions. This angle depends only mildly on the remnant BH spin. The projected amplitude of the fundamental QQNM is always maximized by a viewing angle in a narrow range around $30^{\circ}\lesssim\theta\lesssim 45^{\circ}$. For a non-spinning remnant, the predictions agree on a value of $30^{\circ}$ for the viewing angle. The maximum amplitude viewing angle of the reconstructed simulations all live (within their error bars) in the narrow range given by the predictions.
@@
Figure 9: The magnitude of the decay-corrected amplitudes obtained from spatial mapping of the fundamental quadratic mode with $\tilde{\alpha}=\tilde{\alpha}^{\prime}=(2,2,0,+)$ for simulation 0004. The solid lines indicate the $m=4$ spherical modes, whilst the translucent grey lines indicate odd $m$ modes, and the translucent red lines indicate even $m\neq 4$ modes. The ratios of selected amplitudes are plotted in the lower panels. The values for the structure predicted by second-order PT (expressed in terms of ratios of second-order mode mixing coefficients) are plotted as horizontal dashed lines. The translucent area surrounding the lines on the three panels indicate the NR uncertainty in the values obtained.
In general, error bars are larger for lower remnant spins. This is most clearly demonstrated in the two simulations with the smallest spins, which are considerably harder to map, resulting in a greater sensitivity on the start time used. This effect may be due to higher spin remnants having longer ringdowns, which provides more signal on which to perform the mapping, leading to smaller errors.
Figure 9 shows the reconstructed amplitudes $A^{\ell m}_{\tilde{\alpha}\tilde{\alpha}^{\prime}}$ for the fundamental QQNMs in simulation 0004 as a function of start time. These (decay-corrected) amplitudes show a clear stable behavior across a wide range of start times for several different $\ell$ indices. This is important because, as noted above, in order for these reconstructions to be meaningful it is necessary to detect the feature in multiple spherical harmonics.
@@
Figure 10: The magnitude of the decay-corrected amplitudes obtained from spatial mapping of the QQNM $\tilde{\alpha}=(2,2,0,+)$, $\tilde{\alpha}^{\prime}=(2,-2,0,-)$ for simulation 0004. The solid lines indicate the $m=0$ spherical modes, whilst the translucent grey lines indicate odd $m$ modes, and the translucent red lines indicate even $m\neq 0$ modes. The ratios of selected amplitudes are plotted in the lower panels. The values for the structure predicted by second-order PT (expressed in terms of ratios of second-order mode mixing coefficients) are plotted as horizontal dashed lines. The translucent area surrounding the lines on the three panels indicate the NR uncertainty in the values obtained.
@@
IV.1 $m=0$ modes
Among all possible QQNMs, there is a subset of modes with parent QNMs $\alpha=(\ell,m,n,p)$ and $\alpha^{\prime}=(\ell,-m,n,-p)$. These modes have purely imaginary frequencies and are therefore non-oscillatory, unlike most other QNMs or QQNMs. Because the azimuthal indices of the linear QNMs in the pair sum to zero, the resulting QQNM appears in the azimuthally symmetric $m=0$ sector. Additionally, if $mp>0$, then because both linear QNMs in the pair are prograde and typically excited with large amplitudes. The resulting QQNM is therefore also expected to have a large amplitude. These properties suggest a connection between this subset of QQNMs and the GW memory effect.
To investigate these further, the $\tilde{\alpha}=(2,2,0,+)$, $\tilde{\alpha}^{\prime}=(2,-2,0,-)$ QQNM is spatially mapped, and the amplitudes and their ratios are plotted in Fig. 10. In principle, this combination of prograde parent modes should be the loudest $m=0$ mode in the ringdown, if it is indeed present. The stability of the mapped amplitudes is considerably weaker than in the previous cases, which may indicate that this QQNM is too quiet to be clearly detected. However, flat regions in the amplitudes and upper amplitude ratio give some indication that the mapping is successfully identifying the mode and its corresponding mixing.
@@ V Discussion
This paper has demonstrated that it is possible to determine the specific spatial shape of QNMs and QQNMs. Using CCE waveforms, it has been shown that the angular structure of QNMs can be accurately reconstructed in agreement with first order PT. In analogy with single and multi-mode ringdown fits, the spatial mapping procedure has a mismatch – called the spatial mismatch – which, correspondingly, has a minimum that represents the region around which the QNM is clearly present in the ringdown.
@@
In contrast to first order, second order PT makes no clear prediction about the spatial structure of QQNMs. Hence, spatial mapping presents a method to obtain this structure for a given QQNM and remnant black hole. This is likely to be useful in future precision waveform modelling, where second-order effects will be important.
The spatial mapping has been applied to what is expected to be the loudest QQNM, demonstrating that this method can determine the spatial structure for second order PT. The technique was also applied to the loudest $m=0$ QQNM. This mode is special due to a possible connection with the memory effect. We demonstrate that it may be possible to reconstruct this mode, providing another way to investigate this connection.
While the predictions considered in this work vary, we have shown they are generally qualitatively very similar to each other and the reconstructed spatial structure. Perhaps the most significant finding is that the loudest $\tilde{\alpha}=\tilde{\alpha}^{\prime}=(2,2,0,+)$ QQNM is not visible at $\theta=0$. Instead, the projected amplitude is maximized at a viewing angle of $\theta\approx 35^{\circ}$ for a remnant with spin $\chi_{f}\approx 0.7$. This will hopefully provide a guide for future searches of QQNMs in GW data.
@@ Acknowledgements.
We thank Keefe Mitman, Eliot Finch, Julian Westerweck, and Maximiliano Isi for their very helpful comments on an early draft of this paper. We also thank Keefe Mitman for organizing the public release of the high-fidelity CCE simulations used in this study.
Appendix A Single and multi-mode fitting
This appendix describes the single and multi-mode least-squares fitting used in this paper, carried out using the qnmfits package, in addition to the modifications required to perform spatial mapping. Here, in addition to the abbreviated notation $\alpha=(\ell,m,n,p)$ for the indices of the various QNMs, a similar abbreviated notation $\beta=(\ell,m)$ for the spherical mode indices is also introduced. Recall, spherical $\beta$ indices are placed in superscript while spheroidal $\alpha$ indices are placed in subscript. The perturbation order is suppressed in the notation of this appendix. The strain time series is sampled at discrete times $t_{k}$, with $k=0,1,\ldots,K-1$. The summation convention is not used; all summations are written explicitly.
Starting from the QNM model in Eq. 3 and expanding the spheroidal harmonics in terms of the spherical harmonics using the mode mixing coefficients in Eq. 4, gives the following QNM model for a single spherical harmonic mode,
@@ Appendix A Single and multi-mode fitting
Hereafter, the factors of $r$ (on the left-hand side) and $M$ (on the right-hand side) in Eq. 23 are dropped from our notation for simplicity. When fitting a single spherical harmonic mode, the mode mixing coefficient can be absorbed into a redefinition of the QNM amplitude, $\tilde{C}_{\alpha}=C_{\alpha}\mu^{\beta}_{\alpha}$. The amplitudes $\tilde{C}_{\alpha}$, are determined by minimizing the sum of the squares of the fit residuals,
@@
When fitting a QNM model to multiple spherical harmonic modes simultaneously (as is done in the angle-averaged fits) it is no longer possible to ignore the mode mixing coefficients by a redefinition of the amplitudes. A QNM $(\ell,m,n,p)$ generates a contribution in all the spherical modes $(\ell^{\prime},m^{\prime})$ with $m=m^{\prime}$; this effect is illustrated in the shaded column of plots in Fig. 12. The QNM model in Eq. 23 is now used for all values of $\beta$ for which NR data is available (instead of just a single value) and the QNM amplitudes $C_{\alpha}$ are determined by minimizing the following sum of the squares of the fit residuals summed over all spherical harmonic modes,
@@
Here the vectors and matrices are now formed by stacking together (concatenating) all the spherical harmonic modes. For example, the quantity $\boldsymbol{\mathfrak{h}}$ is now a vector of length $KM$ (where $M$ is the number of spherical harmonics used in the fit; i.e. the number of terms in the sum over $\beta$) formed from stacking all the $K$-vectors $\boldsymbol{\mathfrak{h}}^{\beta}$ together. The order in which the vectors are stacked is arbitrary, but must be done consistently with the construction of the matrix $\boldsymbol{a}$. The quantity $\boldsymbol{a}$ is now a $(KM,N)$ matrix containing the mode mixing and exponential factors in Eq. 26. The quantity $\boldsymbol{C}$ is still a vector of length $N$ containing the QNM amplitudes to be determined. Specifically, in terms of components, these are given by $(\boldsymbol{\mathfrak{h}})_{i}=\mathfrak{h}^{\beta}(t_{k})$ with $\beta=i//K$ and $k=i\%K$, $(\boldsymbol{a})_{ij}=\mu^{\beta}_{j}\exp(-i\omega_{j}(t_{k}-t_{0}))$ again with $\beta=i//K$ and $k=i\%K$, and $\boldsymbol{C}_{j}=C_{j}$. Here, Python notation is used for the modulus and integer division operations. In this form, the numpy.linalg.lstsq function can again be used to efficiently minimize the sum of the squared residuals and find the corresponding QNM amplitudes, $\boldsymbol{C}$.
An angle-averaged fit to multiple spherical harmonic modes is illustrated in Fig. 12 for simulation 0013. Each panel shows both the data, $\mathfrak{h}^{\beta}$, and the QNM model, $h^{\beta}$, for one spherical harmonic mode. The model gives a good fit to all the modes, as can be seen by examining the residuals. It should be stressed that each panel in this figure is not a separate fit; they are all part of the same fit. This is necessary because of the mode mixing effect and is illustrated graphically by highlighting the effect of a single QNM (the fundamental $\alpha=(2,2,0,+)$ mode) in different spherical harmonics (see the three plots with $m=2$ in the highlighted column). In this example, because mode mixing only occurs between modes with the same index $m$, the different columns of this figure are, in fact, independent. However, when mapping (see Sec. II.3) the shape of a particular mode, the mode mixing coefficients are effectively promoted to free parameters and all parts of this plot would be coupled.
The angle-averaged fitting procedure can be extended to include a spatially mapped mode, $\tilde{\alpha}$. Using the mapping model from Eq. 11 (see also Eq. 12), the sum of the squared residuals in Eq. 26 now becomes,
@@
Figure 12: Illustration of an angle-averaged multimode QNM fit to NR data. Shown in blue is the NR waveform $\mathfrak{h}^{\beta}(t)$ for the CCE 0013 simulation which, because it comes from a precessing asymmetric binary (see Table 1), contains significant power spread across many spherical harmonic modes. Here, for the purpose of illustration, modes up to and including $\ell_{\rm max}=4$ are plotted (real [solid] and imaginary [dotted] parts shown separately) arranged in a grid with $m$ and $\ell$ increasing along the rows and down the columns respectively. Note that the $y$-axis range is shared for all panels in a given row (i.e. with the same $\ell$) but changes down the columns; the time range on the $x$-axis is the same in all panels. Shown in green is a QNM model angle-averaged fit to the NR data. The QNM model, $h^{\beta}(t)$, starts at $t_{0}=10M$ and includes all QNMs (prograde and retrograde) up to and including those with $\ell_{\rm max}=4$ and $n_{\rm max}=1$; the total number of damped sinusoids in this fit is 84. The residuals (data - model) are plotted in black on each panel, shifted vertically down, and multiplied by a factor of 10 for clarity. It should be emphasized that the angle-averaged fit is not separately fitting a sum of damped sinusoids to each spherical mode; rather the intrinsic amplitude and phase of each QNM, $C_{\alpha}$, is determined by fitting to all the spherical harmonic modes simultaneously with its contribution in each multiplied by the mode mixing coefficient $\mu^{\beta^{\prime}}_{\alpha}$. Only the amplitude $C_{\alpha}$ is free to vary in the fit, the mode mixing coefficients $\mu^{\beta^{\prime}}_{\alpha}$ are determined by the final spin of the remnant. Because, as a consequence of axial symmetry, the mode mixing coefficients satisfy $\mu^{\beta^{\prime}}_{\alpha}\propto\delta_{m}^{m^{\prime}}$ the mixing occurs only within the columns of this figure. In order to emphasize the role of mode mixing, the contribution of the fundamental QNM with $\alpha=(\ell,m,n,p)=(2,2,0,+)$ is isolated and plotted in orange (shifted vertically up for clarity) in the highlighted $m=2$ column. The formula for the amplitude of this contribution is indicated in each panel by the inset legend. This fundamental QNM dominates the spherical $\ell=m=2$ mode signal, but is also significant in the $\ell=3$, $m=2$ mode, with smaller contributions for higher $\ell$ values. It should be noted that the remnant BH in simulation 0013 has a small final spin of $\chi_{f}=0.43$ which means that the mode mixing is less prominent here than in other cases.
Appendix B Simultaneous mapping of multiple QNMs

### v2

@@ IV Quadratic QNM results
In principle, second-order PT is capable of making an unambiguous prediction for the second-order part of the metric perturbation, including all of the QQNMs. While work in this direction has been attempted [30, 59], there is no analog of the spheroidal harmonics for QQNMs, i.e. there does not exist a simple, universal family of functions that describe the shape of these features for all perturbed BHs.
In the absence of complete calculations at second order, there have been a number of guesses, or predictions, made in the literature for the spatial shape of the QQNMs. These are inspired by second-order PT. Because the QQNM is sourced by two linear QNMs, some of these assume that the shape $F_{\alpha\alpha^{\prime}}$ (see Eq. 5) will be given by a quadratic combination of spheroidal harmonics. In this section, several of these incomplete second-order PT predictions are outlined and compared to the spatially reconstructed modes.
Prediction A: A naive guess for these functions is to take the product of two parent spheroidal harmonics, ${}_{-2}S_{\alpha}(\theta,\phi)$ and ${}_{-2}S_{\alpha^{\prime}}(\theta,\phi)$. Then, the spatial shape of the QQNM takes the form
@@
However, while this gives qualitatively the correct dependence on the azimuthal $\phi$ coordinate, it gives the wrong $\theta$ dependence (see, for example, the reconstruction of the dominant QQNM in Fig. 7). In particular, it predicts that the dominant QQNM’s largest observed amplitude would be at the north pole ($\theta=0$). This is in disagreement with the reconstruction, which shows a finite angle $\theta_{\rm max}\approx 35^{\circ}$ for this maximum. This is further quantified in Fig. 8, where $\theta_{\mathrm{max}}$ is plotted for the $\alpha=\alpha^{\prime}=(2,2,0,+)$ QQNM for all simulations.
This prediction cannot be correct because the resulting $F_{\alpha\alpha^{\prime}}$ does not have the required -2 spin weight for a GW field. Instead, $F_{\alpha\alpha^{\prime}}$ in Eq. 18 has a spin weight of -4 resulting from the fact that the product of a spin-weight $s$ function and a spin weight $s^{\prime}$ function has spin-weight $s+s^{\prime}$.
Prediction B: This issue can be resolved by using a product of spheroidal harmonics with spin weights that sum to -2. There is not a unique choice for this. The possibilities for the fundamental QQNM with $\alpha=\alpha^{\prime}=(2,2,0,+)$ are
@@
Prediction C: There is another proposal given in [42]. This assumes that the shape of the QQNM is given by a single spheroidal harmonic which, rather than being defined by its indices, has a modified spheroidicity defined using the frequency $\omega_{\alpha}+\omega_{\alpha^{\prime}}$ of the QQNM,
$\mathbf{C:}\quad F_{\alpha\alpha^{\prime}}(\theta,\phi)\propto{}_{-2}S_{\lambda\mu}\big{(}i\chi_{f}\left[\omega_{\alpha}+\omega_{\alpha^{\prime}}\right],\theta,\phi\big{)},$ (21)
where $\lambda=\ell+\ell^{\prime}$ and $\mu=m+m^{\prime}$ and, for clarity, we have temporarily reverted to the original notation for the spheroidal harmonics introduced in Eq. 2. Here, the notation for the spheroidal harmonics from Ref. [43] is used.
Prediction D: An alternative resolution to the spin-weight issue is to take the product of two parent spheroidal harmonics with spin weights that do not sum to $-2$, but act on them with the appropriate spin-weight raising or lowering operator to obtain the required spin-weight $-2$ function [60]. For example, taking the product of two spin-weight $-2$ spheroidal harmonics as in prediction A but acting on them twice using the spin-weight raising operator gives,
@@
For each of these predictions, a corresponding quadratic mode-mixing coefficient $\mu^{\ell m}_{\alpha\alpha^{\prime}}$ can be computed. These three predictions are compared in Fig. 7 against the reconstruction of the fundamental QQNM in simulation 0001. Prediction A is clearly disfavored and is not discussed further. Visually, there is little difference between the other predictions.
Interestingly, the B, C and D predictions all correctly suggest that the projected amplitude of the fundamental QQNM is maximized at a nonzero viewing angle, away from the axis of the remnant BH’s final spin. This has important consequences for the observability of this QQNM in GW data. This angle is investigated further in Fig. 8 where it is plotted as a function of the spin magnitude of the remnant for all remaining predictions. This angle depends only mildly on the remnant BH spin. The projected amplitude of the fundamental QQNM is always maximized by a viewing angle in a narrow range around $30^{\circ}\lesssim\theta\lesssim 45^{\circ}$. For a nonspinning remnant, the predictions reduce to a value of $30^{\circ}$, corresponding to the maximum amplitude angle of the ${}_{-2}Y^{44}$ spherical harmonic. As spin increases, the contribution of higher $\ell$ spherical modes becomes increasingly significant, resulting in the unique structures identified in the figure. The maximum amplitude viewing angles of the reconstructed simulations all live (within their error bars) in the narrow range given by the predictions.
@@
Figure 9: The magnitude of the decay-corrected amplitudes obtained from spatial mapping of the fundamental quadratic mode with $\tilde{\alpha}=\tilde{\alpha}^{\prime}=(2,2,0,+)$ for simulation 0004. The solid lines indicate the $m=4$ spherical modes, whilst the translucent grey lines indicate odd $m$ modes, and the translucent red lines indicate even $m\neq 4$ modes. The ratios of selected amplitudes are plotted in the lower panels. The values for the structure predicted by second-order PT (expressed in terms of ratios of second-order mode-mixing coefficients) are plotted as horizontal dashed lines. The translucent areas surrounding the lines on the three panels indicate the NR uncertainty in the values obtained.
In general, error bars are larger for lower remnant spins. This is most clearly demonstrated in the two simulations with the smallest spins, which are considerably harder to map, resulting in a greater sensitivity to the start time used. This effect may be due to higher spin remnants having longer ringdowns, which provides more signal on which to perform the mapping, leading to smaller errors.
Figure 9 shows the reconstructed amplitudes $A^{\ell m}_{\tilde{\alpha}\tilde{\alpha}^{\prime}}$ for the fundamental QQNMs in simulation 0004 as a function of the start time. These (decay-corrected) amplitudes show a clear stable behavior across a wide range of start times for several different $\ell$ indices. This is important because, as noted above, in order for these reconstructions to be meaningful, it is necessary to detect the feature in multiple spherical harmonics.
@@
Figure 10: The magnitude of the decay-corrected amplitudes obtained from spatial mapping of the QQNM $\tilde{\alpha}=(2,2,0,+)$ and $\tilde{\alpha}^{\prime}=(2,-2,0,-)$ for simulation 0004. The solid lines indicate the $m=0$ spherical modes, whilst the translucent grey lines indicate odd $m$ modes, and the translucent red lines indicate even $m\neq 0$ modes. The ratios of selected amplitudes are plotted in the lower panels. The values for the structure predicted by second-order PT (expressed in terms of ratios of second-order mode-mixing coefficients) are plotted as horizontal dashed lines. The translucent areas surrounding the lines on the three panels indicate the NR uncertainty in the values obtained.
@@
IV.1 The $m=0$ modes
Among all possible QQNMs, there is a subset of modes with parent QNMs $\alpha=(\ell,m,n,p)$ and $\alpha^{\prime}=(\ell,-m,n,-p)$. These modes have purely imaginary frequencies and are, therefore, non-oscillatory, unlike most other QNMs or QQNMs. Because the azimuthal indices of the linear QNMs in the pair sum to zero, the resulting QQNM appears in the azimuthally symmetric $m=0$ sector. Additionally, if $mp>0$, then because both linear QNMs in the pair are prograde and typically excited with large amplitudes, the resulting QQNM is, therefore, also expected to have a large amplitude. These properties suggest a connection between this subset of QQNMs and the GW memory effect.
To investigate these further, the $\tilde{\alpha}=(2,2,0,+)$, $\tilde{\alpha}^{\prime}=(2,-2,0,-)$ QQNM is spatially mapped, and the amplitudes and their ratios are plotted in Fig. 10. In principle, this combination of prograde parent modes should be the loudest $m=0$ mode in the ringdown, if it is indeed present. The stability of the mapped amplitudes is considerably weaker than in the previous cases, which may indicate that this QQNM is too quiet to be clearly detected. However, flat regions in the amplitudes and the upper amplitude ratio give some indication that the mapping is successfully identifying the mode and its corresponding mixing.
@@ V Discussion
This paper has demonstrated that it is possible to determine the specific spatial shape of QNMs and QQNMs. Using CCE waveforms, it has been shown that the angular structure of QNMs can be accurately reconstructed in agreement with first-order PT. In analogy with single- and multi-mode ringdown fits, the spatial mapping procedure has a mismatch – called the spatial mismatch – which, correspondingly, has a minimum that represents the region around which the QNM is clearly present in the ringdown.
@@
In contrast to first order, second-order PT makes no clear prediction about the spatial structure of QQNMs. Hence, spatial mapping presents a method to obtain this structure for a given QQNM and remnant black hole. This is likely to be useful in future precision waveform modeling, where second-order effects will be important.
The spatial mapping has been applied to what is expected to be the loudest QQNM, demonstrating that this method can determine the spatial structure for second-order PT. The technique was also applied to the loudest $m=0$ QQNM. This mode is special due to a possible connection with the memory effect. We demonstrated that it may be possible to reconstruct this mode, providing another way to investigate this connection.
While the predictions considered in this work varied, we have shown that they were generally qualitatively very similar to each other and the reconstructed spatial structure. Perhaps the most significant finding is that the loudest $\tilde{\alpha}=\tilde{\alpha}^{\prime}=(2,2,0,+)$ QQNM was not visible at $\theta=0$. Instead, the projected amplitude was maximized at a viewing angle of $\theta\approx 35^{\circ}$ for a remnant with spin $\chi_{f}\approx 0.7$. This will hopefully provide a guide for future searches of QQNMs in GW data.
@@ Acknowledgements.
We thank Keefe Mitman, Eliot Finch, Julian Westerweck, and Maximiliano Isi for their helpful comments on a draft of this paper. We also thank the anonymous referee for their very helpful report. Finally, we thank Keefe Mitman for organizing the public release of the high-fidelity CCE simulations used in this study.
Appendix A Single- and multi-mode fitting
This appendix describes the single- and multi-mode least-squares fitting used in this paper, carried out using the qnmfits package, in addition to the modifications required to perform spatial mapping. Here, in addition to the abbreviated notation $\alpha=(\ell,m,n,p)$ for the indices of the various QNMs, a similar abbreviated notation $\beta=(\ell,m)$ for the spherical-mode indices is also introduced. Recall, spherical $\beta$ indices are placed in superscript while spheroidal $\alpha$ indices are placed in subscript. The perturbation order is suppressed in the notation of this appendix. The strain time series is sampled at discrete times $t_{k}$, with $k=0,1,\ldots,K-1$. The summation convention is not used; all summations are written explicitly.
Starting from the QNM model in Eq. 3 and expanding the spheroidal harmonics in terms of the spherical harmonics using the mode-mixing coefficients in Eq. 4, gives the following QNM model for a single spherical harmonic mode:
@@ Appendix A Single- and multi-mode fitting
Hereafter, the factors of $r$ (on the left-hand side) and $M$ (on the right-hand side) in Eq. 23 are dropped from our notation for simplicity. When fitting a single spherical harmonic mode, the mode-mixing coefficient can be absorbed into a redefinition of the QNM amplitude, $\tilde{C}_{\alpha}=C_{\alpha}\mu^{\beta}_{\alpha}$. The amplitudes $\tilde{C}_{\alpha}$, are determined by minimizing the sum of the squares of the fit residuals,
@@
When fitting a QNM model to multiple spherical harmonic modes simultaneously (as is done in the angle-averaged fits) it is no longer possible to ignore the mode-mixing coefficients by a redefinition of the amplitudes. A QNM $(\ell,m,n,p)$ generates a contribution in all the spherical modes $(\ell^{\prime},m^{\prime})$ with $m=m^{\prime}$, and this effect is illustrated in the shaded column of plots in Fig. 12. The QNM model in Eq. 23 is now used for all values of $\beta$ for which NR data are available (instead of just a single value), and the QNM amplitudes $C_{\alpha}$ are determined by minimizing the following sum of the squares of the fit residuals summed over all spherical harmonic modes:
@@
Here, the vectors and matrices are now formed by stacking together (concatenating) all the spherical harmonic modes. For example, the quantity $\boldsymbol{\mathfrak{h}}$ is now a vector of length $KM$ (where $M$ is the number of spherical harmonics used in the fit, i.e.  the number of terms in the sum over $\beta$) formed from stacking all the $K$ vectors $\boldsymbol{\mathfrak{h}}^{\beta}$ together. The order in which the vectors are stacked is arbitrary, but must be done consistently with the construction of the matrix $\boldsymbol{a}$. The quantity $\boldsymbol{a}$ is now a $(KM,N)$ matrix containing the mode-mixing and exponential factors in Eq. 26. The quantity $\boldsymbol{C}$ is still a vector of length $N$ containing the QNM amplitudes to be determined. Specifically, in terms of components, these are given by $(\boldsymbol{\mathfrak{h}})_{i}=\mathfrak{h}^{\beta}(t_{k})$ with $\beta=i//K$ and $k=i\%K$, $(\boldsymbol{a})_{ij}=\mu^{\beta}_{j}\exp(-i\omega_{j}(t_{k}-t_{0}))$ again with $\beta=i//K$ and $k=i\%K$, and $\boldsymbol{C}_{j}=C_{j}$. Here, Python notation is used for the modulus and integer division operations. In this form, the numpy.linalg.lstsq function can again be used to efficiently minimize the sum of the squared residuals and find the corresponding QNM amplitudes, $\boldsymbol{C}$.
An angle-averaged fit to multiple spherical harmonic modes is illustrated in Fig. 12 for simulation 0013. Each panel shows both the data, $\mathfrak{h}^{\beta}$, and the QNM model, $h^{\beta}$, for one spherical harmonic mode. The model gives a good fit to all the modes, as can be seen by examining the residuals. It should be stressed that each panel in this figure is not a separate fit – they are all part of the same fit. This is necessary because of the mode-mixing effect and is illustrated graphically by highlighting the effect of a single QNM (the fundamental $\alpha=(2,2,0,+)$ mode) in different spherical harmonics (see the three plots with $m=2$ in the highlighted column). In this example, because mode mixing only occurs between modes with the same index $m$, the different columns of this figure are, in fact, independent. However, when mapping (see Sec. II.3) the shape of a particular mode, the mode-mixing coefficients are effectively promoted to free parameters and all parts of this plot would be coupled.
The angle-averaged fitting procedure can be extended to include a spatially mapped mode, $\tilde{\alpha}$. Using the mapping model from Eq. 11 (see also Eq. 12), the sum of the squared residuals in Eq. 26 now becomes
@@
Figure 12: Illustration of an angle-averaged multimode QNM fit to NR data. Shown in blue is the NR waveform $\mathfrak{h}^{\beta}(t)$ for the CCE 0013 simulation, which, because it comes from a precessing asymmetric binary (see Table 1), contains significant power spread across many spherical harmonic modes. Here, for the purpose of illustration, modes up to and including $\ell_{\rm max}=4$ are plotted (real [solid] and imaginary [dotted] parts shown separately), arranged in a grid with $m$ and $\ell$ increasing along the rows and down the columns, respectively. Note that the $y$-axis range is shared for all panels in a given row (i.e.  with the same $\ell$) but changes down the columns, while the time range on the $x$-axis is the same in all panels. Shown in green is a QNM model angle-averaged fit to the NR data. The QNM model, $h^{\beta}(t)$, starts at $t_{0}=10M$ and includes all QNMs (prograde and retrograde) up to and including those with $\ell_{\rm max}=4$ and $n_{\rm max}=1$. The total number of damped sinusoids in this fit is 84. The residuals (data minus model) are plotted in black on each panel, shifted vertically down, and multiplied by a factor of 10 for clarity. It should be emphasized that the angle-averaged fit is not separately fitting a sum of damped sinusoids to each spherical mode; rather, the intrinsic amplitude and phase of each QNM, $C_{\alpha}$, is determined by fitting to all the spherical harmonic modes simultaneously, with its contribution in each multiplied by the mode-mixing coefficient $\mu^{\beta^{\prime}}_{\alpha}$. Only the amplitude $C_{\alpha}$ is free to vary in the fit – the mode-mixing coefficients $\mu^{\beta^{\prime}}_{\alpha}$ are determined by the final spin of the remnant. Because, as a consequence of axial symmetry, the mode-mixing coefficients satisfy $\mu^{\beta^{\prime}}_{\alpha}\propto\delta_{m}^{m^{\prime}}$, the mixing occurs only within the columns of this figure. In order to emphasize the role of mode mixing, the contribution of the fundamental QNM with $\alpha=(\ell,m,n,p)=(2,2,0,+)$ is isolated and plotted in orange (shifted vertically up for clarity) in the highlighted $m=2$ column. The formula for the amplitude of this contribution is indicated in each panel by the inset legend. This fundamental QNM dominates the spherical $\ell=m=2$ mode signal, but is also significant in the $\ell=3$, $m=2$ mode, with smaller contributions for higher $\ell$ values. It should be noted that the remnant BH in simulation 0013 has a small final spin of $\chi_{f}=0.43$, which means that the mode mixing is less prominent here than in other cases.
Appendix B Effect of varying the number of overtones in a QNM reconstruction
To address the sensitivity of spatial mapping to the choice of QNM content in the model, the figures in this paper can be reproduced for varying values of $n_{\rm max}$ and $\ell_{\rm max}$. An example of this is shown in Fig. 13, where the left panel of Fig. 3 has been reproduced for $n_{\rm max}\in\{4,5,6\}$. The resulting figures are extremely similar to the $n_{\rm max}=7$ case, highlighting the stability of the “kitchen sink” approach to changes in the choice of $n_{\rm max}$.
Figure 13: The effect of changing the overtone content in spatial mapping is demonstrated by reducing $n_{\rm max}$ to 4, 5, or 6. The technique can be seen to be robust against these changes, supporting the “kitchen sink” approach to this work. These figures should be compared with the left panel of Fig. 3, which was produced using $n_{\rm max}=7$.
Appendix C Simultaneous mapping of multiple QNMs